Frank is 5 years younger than Lawrence, four years later, lawrence will be twice as old as frank find their present age
Answer by lwsshak3(11628)

Frank is 5 years younger than Lawrence, four years later, lawrence will be twice as old as frank find their present age.
***
let x=Lawrence's age
x-5=Frank's age
..
4 yrs later:
x+4=Lawrence's age
x-5+4=x-1=Frank's age
x+4=2(x-1)
x+4=2x-2
x=6
x-5=1
Lawrence's age=6
Frank's age=1